OpenStudy (anonymous):

Mr. Crosson selects two students from his class to present their reports on Spain. He places the names of each student in a box and picks one at a time. The class has 8 boys and 10 girls. What is the probability that Mr. Crosson will pick two boys in a row if he does not put the first name back in the box? Round your answer to the nearest percent. A.16% B.18% C.17% D.19%

OpenStudy (anonymous):

\(\frac{8}{18}\times \frac{7}{17}\)

OpenStudy (anonymous):

so that would make it 18%

OpenStudy (anonymous):

?